OpenStudy (anonymous):

Please check my answers Use the quadratic formula to solve the equation. If necessary, round to the nearest hundredth. 2a^2-46a+252=0 9, 14 - my answer –9, –14 –18, 28 18, 28 2) The table shows the estimated number of deer living in a forest over a five-year period. Are the data best represented by a linear, exponential, or quadratic model? Write an equation to model the data. Year Estimated Population 0 91 1 70 2 54 3 42 4 32

OpenStudy (dinamix):

yup its true 9 and 14

OpenStudy (anonymous):

A) exponential; y = 91*0.77^x - my answer B) linear; y = 0.77x+91 C) quadratic; y = 91x^2+0.77 D) quadratic; y = 0.77x^2+91

OpenStudy (anonymous):

Use any method to solve the equation. If necessary, round to the nearest hundredth. 9x^2 = 17 3, –3 - my answer 0.73, –0.73 4.12, –4.12 1.37, –1.37

OpenStudy (anonymous):

Find the number of real number solutions for the equation. x^2+5x+7-=0 1 2 Can not be determined from the graph 0 - my answer Graph the set of points. Which model is most appropriate for the set? (–6, –1), (–3, 2), (–1, 4), (2, 7) exponential quadratic linear - my answer

OpenStudy (anonymous):

Solve the equation by completing the square. Round to the nearest hundredth if necessary. x^2-4x=5 1.73, –1.73 5, –1 - my answer 11, –7 1, 3 Use any method to solve the equation. If necessary, round to the nearest hundredth. 0.85, –0.85 - my answer 2.83, –2.83 3.32, –3.32 1.17, –1.17

OpenStudy (anonymous):

A) exponential; y = 91*0.77^x - my answer B) linear; y = 0.77x+91 C) quadratic; y = 91x^2+0.77 D) quadratic; y = 0.77x^2+91 correct Use any method to solve the equation. If necessary, round to the nearest hundredth. 9x^2 = 17 3, –3 - my answer 0.73, –0.73 4.12, –4.12 1.37, –1.37 i think it is the last one

OpenStudy (dinamix):

x(1) = -5/2 -i((3)^1/2)/2 and x(2) = -5/2+i((3)^1/2)/2

OpenStudy (dinamix):

@ this solution for x^2+4x+7 =o @proud_yemeniah_

OpenStudy (anonymous):

is my answer wrong @dinamix

OpenStudy (dinamix):

Can not be determined from the graph i think

OpenStudy (anonymous):

okay thanks can u check on the questions I have left

OpenStudy (dinamix):

@proud_yemeniah_ lniear is correct

OpenStudy (anonymous):

can u help me with whats left @dinamix Solve the equation by completing the square. Round to the nearest hundredth if necessary. x^2-4x=5 1.73, –1.73 5, –1 - my answer 11, –7 1, 3 Use any method to solve the equation. If necessary, round to the nearest hundredth. 11x^2 = 8 0.85, –0.85 - my answer 2.83, –2.83 3.32, –3.32 1.17, –1.17

OpenStudy (dinamix):

5,-1 is true

OpenStudy (dinamix):

0.85;-0.85 is true u are really cool

OpenStudy (anonymous):

Thank you both! @dinamix @ASAAD123